OpenStudy (anonymous):

PLEASE HELP DO NOT UNDERSTAND THANKS Find a polynomial function f(x) of the indicated degree with integer coefficients that possesses the given rational zeros degree=4; zeros= -3, 1/2,1,2

OpenStudy (blake57roger):

Let a, b, c be the roots. factored form: f(x)=k(x-a)(x-b)(x-c) with k being a constant. You can add extra roots (like p) by tacking an extra factor (like (x-p) ) to the end. You can manipulate the degree by raising one of the roots to an exponent. .

OpenStudy (anonymous):

so k would be 4 in this case correct?

OpenStudy (amorfide):

degree 4 means there are 4 roots to the function there fore you want to end up with x^4 K is just a constant as blake said, blake also said you can add an extra factor (x-p) which would give f(x)=k(x-a)(x-b)(x-c)(x-p) or you can have a repeated root which is his point of manipulating